Session: Feb 2000 (session 1)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Morsiani at Sun Feb 27 14:33:44 MET 2000) Minor failures
Lost 1 and 2 scan Start 20 seconds later in 4,9,10 and 11 scan

Torun (K.Borkowski at Sun Feb 27 14:59:26 MET 2000) Success
No known problems.

Medicina (morsiani at Mon Feb 28 09:43:18 MET 2000) Minor failures
the scans lost are due to operator error not at setup problems. The time needed to change rx must be slightly increased

Effelsberg (A. Kraus at Mon Feb 28 12:26:42 MET 2000) Success

Medicina (A. Polatidis at Tue Feb 29 08:43:36 MET 2000) Success
Observed 8 GHz RCP only.

Westerbork (Tony Foley at Tue Feb 29 14:03:31 MET 2000) Minor failures
LCP and RCP are swapped in the 6cm scans; SORRY

Jodrell (Alastair Gunn at Thu Mar 2 05:34:29 MET 2000) Minor failures
Telescope parked due to high winds at start of experiment. On source at about 11:53 UT. No other known problems.

Noto (Noto VLBI friend at Thu Mar 2 12:06:45 MET 2000) Success

Calibration (Alastair Gunn at Wed Apr 19 13:08:45 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/gah03.antab for details